OpenStudy (shealah):

The second of three numbers is 1 less than the first. The third number is 11 less than the second. If the frst number is twice as large as the the third, find the three numbers.

zepdrix (zepdrix):

oo system of equations! these are fun :)

zepdrix (zepdrix):

So here is our system :o (f=first) (s=second) (t=third) s=f-1 <-- second number 1 less than the first t=s-11 <-- third number 11 less than then second f=2t <-- the first number is twice the third number So three equations.. hmmmm what can we do :)

zepdrix (zepdrix):

Either substitution or elimination will work, are you comfortable with either method? :D

OpenStudy (anonymous):

i am not sure i would write three equations, although that would also work i might call the second number say \(x\) then the first number is \(x+1\) and the third is \(x-11\) and you know also that the first number \(x+1\) is equal to \(2(x-11)\), twice the third number so you can write \[x+1=2(x-11)\] and solve for \(x\)

OpenStudy (shealah):

Thanks :)